# SCMIL: Sparse Context-aware Multiple Instance Learning for Predicting Cancer Survival Probability Distribution in Whole Slide Images

### v1

@@ 2.2 Sparse Context-aware Self-Attention (SCSA)
$\alpha_{i}=\frac{exp(a^{T}(tanh(VH_{i}^{\prime T})\odot\sigma(UH_{i}^{\prime T})))}{\sum_{k=1}^{n}exp(w^{T}(tanh(VH_{k}^{\prime T})\odot\sigma(UH_{k}^{\prime T})))}$ (4)
@@ 2.3 RegisterMDN
The Mixed Density Network (MDN) translates the input to a probability distribution. We adopt Gaussian distributions as the components of the MDN, assuming that the number of components is $K$. We utilize the WSI-level features $Feat^{\prime}$, the mean vector $P_{m}$, and the standard deviation vector $P_{v}$ as the input of our RegisterMDN. Both $P_{m}$ and $P_{v}$ are learnable parameters and learn the survival risk characteristics of the specific cancer during the training phase. $Feat^{\prime}$, $P_{m}$, and $P_{var}$ through the neural networks to produce the weights $\lambda_{i}(Feat^{\prime})$, means $\mu_{i}(P_{m})$, and variances $\sigma_{i}(P_{v})$ of the mixture model. Consequently, we can get the Probability Density Function (PDF):
$PDF(y|Feat^{\prime},P_{m},P_{v})=\sum_{i=1}^{K}\lambda_{i}(Feat^{\prime})\mathcal{N}(y|\mu_{i}(P_{m}),\sigma_{i}(P_{v}))$ (6)
@@
$DPDF(t|Feat^{\prime},P_{m},P_{v})=|\frac{\mathrm{d}g^{-1}}{\mathrm{d}t}|\sum_{i=1]}^{K}\lambda_{i}(Feat^{\prime})\mathcal{N}(g^{-1}(t)|\mu_{i}(P_{m}),\sigma_{i}(P_{v}))$ (7)

### v2

@@ 2.2 Sparse Context-aware Self-Attention (SCSA)
$\alpha_{i}=\frac{exp(a^{T}(tanh(VH_{i}^{\prime T})\odot\sigma(UH_{i}^{\prime T})))}{\sum_{k=1}^{n}exp(a^{T}(tanh(VH_{k}^{\prime T})\odot\sigma(UH_{k}^{\prime T})))}$ (4)
@@ 2.3 RegisterMDN
The Mixed Density Network (MDN) translates the input to a probability distribution. We adopt Gaussian distributions as the components of the MDN, assuming that the number of components is $K$. We utilize the WSI-level features $Feat^{\prime}$, the mean vector $P_{m}$, and the standard deviation vector $P_{v}$ as the input of our RegisterMDN. Both $P_{m}$ and $P_{v}$ are learnable parameters and learn the survival risk characteristics of the specific cancer during the training phase. $Feat^{\prime}$, $P_{m}$, and $P_{var}$ through the neural networks to produce the weights $\lambda_{i}(Feat^{\prime})$, means $\mu_{i}(P_{m})$, and variances $\sigma_{i}^{2}(P_{v})$ of the mixture model. Consequently, we can get the Probability Density Function (PDF):
$PDF(y|Feat^{\prime},P_{m},P_{v})=\sum_{i=1}^{K}\lambda_{i}(Feat^{\prime})\mathcal{N}(y|\mu_{i}(P_{m}),\sigma_{i}^{2}(P_{v}))$ (6)
@@
$DPDF(t|Feat^{\prime},P_{m},P_{v})=|\frac{\mathrm{d}g^{-1}}{\mathrm{d}t}|\sum_{i=1]}^{K}\lambda_{i}(Feat^{\prime})\mathcal{N}(g^{-1}(t)|\mu_{i}(P_{m}),\sigma_{i}^{2}(P_{v}))$ (7)